OpenStudy (anonymous):

A group of roommates decide to split the cost of a new sofa. If s is the cost of the sofa and n is the number of roommates, the formula c= s/n determines each person's share of the cost. If the new sofa costs $955 and 4 roommates live in the house, what is each person's share? A. $382.00 B. $238.75 C. $221.25 D. $95.50

OpenStudy (kamibug):

Plug the known info into the formula. s = 955 and n = 4 c = s/n c = 955/4 Just divide to get your answer! :)